OpenStudy (anonymous):

Factor and Solve x^2 - 15x + 50=0

OpenStudy (anonymous):

please hurry. Almost out of time

OpenStudy (triciaal):

what are the factors of 50 that add up to -15?

OpenStudy (anonymous):

(x-10)(x-5)=0 x=10 x=5

OpenStudy (anonymous):

10, -5

OpenStudy (triciaal):

no the signs are important

OpenStudy (anonymous):

the value of x are 10 and 5.

OpenStudy (triciaal):

positive 50 so both -ve or both positive

OpenStudy (anonymous):

(x-10)(x-5)=0 that's the factor. x=10 x=5

OpenStudy (triciaal):

you have -15 so both factors are -ve

OpenStudy (triciaal):

-10 -5 = -15 (-10)(-5) = 50

OpenStudy (triciaal):

(x-10)(x-5)

OpenStudy (triciaal):

to solve x - 10 = 0 or x-5 = 0 x = 10 ; x = 5

OpenStudy (anonymous):

Factor and Solve x^2 - 15x + 50=0 (x-10) (x-5) =0 By distributing above you have the answer of x^2 - 15x + 50=0. Thus, x-10=0 and x-5=0 . The value of x are 10 and 5 to solve the equation.

OpenStudy (anonymous):

ok that's it end.